OpenStudy (anonymous):

Determine iterates of f(x) = 1.x(1 – x)x0 = 0.8

OpenStudy (welshfella):

f(0) = 1.5(0.8)(1 - 0.8)

OpenStudy (anonymous):

Is there a formula for this?

OpenStudy (welshfella):

i should have written f(0.8) not f(0)

OpenStudy (anonymous):

Mhmm, it's a lot easier when simplified.

OpenStudy (welshfella):

calculate the value for f(0.8)

OpenStudy (anonymous):

How do I do that?

OpenStudy (welshfella):

yes we could simplify first 1.5x(1 - x0 = 1.5x - 1.5x^2 f(0.8) = 1.5*0.8 - 1.5*0.8^2

OpenStudy (welshfella):

sorry gotta go right now

OpenStudy (anonymous):

Thanks anyway.